# CyberRAG: An agentic RAG cyber attack classification and reporting tool

### v1

@@ 4.1 Knowledge Bases
For the purposes of this work, we constructed three distinct knowledge bases, one for each type of web-based attack under investigation The system is designed to distinguish among three specific attack types: SSTI, SQL Injection, and Cross-Site Scripting (XSS). To achieve a deep understanding of these attack vectors, dedicated and categorized documents are required.
@@ 4.2 Attacks’ Dataset
For the Server-Side Template Injection (SSTI) category, no ready-to-use dataset is available. Consequently, we generated this dataset manually using curated payloads from the PayloadAllTheThings SSTI section in conjunction with outputs from the SSTImap tool. The positive class consists of diverse payloads known to trigger SSTI vulnerabilities across multiple template engines. In contrast, the negative class comprises a variety of benign strings such as mail, names or numbers that do not result in SSTI behavior, serving as clean input examples. This dataset can be download by password at [29].
@@ 5 Experimental Results
More in depth, each individual component of CyberRAG has been validated, and an assessment of the performance and reliability of the whole system has been performed. This includes the evaluation of the accuracy of the LLM-based classifiers, the effectiveness of the RAG-based explanation module, and the coherence and usability of the final outputs generated by the complete agentic RAG pipeline.
@@ 5.1 LLM-based Classifier for Attack Identification
Figure 3 shows the classification performance of the BERT-based classifiers, used to build the classification tool, when trained independently on each attack type. All models (bert-base-uncased, albert-base-v2, distilbert-base-uncased, roberta-base) were trained and tested for each attack class considered. All classifiers taken into account have been trained with identical hyperparameters: a maximum of 30 epochs and a batch size of 32. The optimizer is RectifiedAdam with a learning rate of $3\times 10^{-5}$. The loss function is binary cross-entropy.
The classifiers were evaluated by using 6 different metrics, three of which are based on the computation of the Area Under the Curve (AUC) for the precision-recall (PR) and receiving operating characteristic (ROC). The other three are the precision (Prec), the F1-score and the binary accuracy (Bin Acc).
@@ 5.3 Evaluation of Generated Explanations
To evaluate the quality of contextualized explanations generated by our RAG framework, we conducted an extensive benchmark across five different open-weight language models, all with approximately 7 billion parameters. The selected models are: DeepSeek-R1 7B [33], Gemma 3B:4B [34], LLaMA3.1 8B [35], Mistral 7B [36], and Qwen2.5 7B [37]. All of these models were made available and executed within the Ollama environment [38], enabling unified access and deployment for comparative evaluation.
@@ 5.3.2 Evaluation via LLM-as-a-Judge
    General Explanation Evaluation: For each generated explanation, GPT-4 rated it on a 1–5 scale for clarity, informativeness, and semantic alignment with the input payload and the attack label.
@@
Table 3: GPT-4 based LLM-judge scores (1–5 scale) per model.
@@ 5.5 Robustness Evaluation
[…]
@@ 6 Discussion
The usage of multiple classifiers allows for a modular architecture that contributes to the robustness of the system. If one classifier performs poorly due to data imbalance or ambiguity in one attack category, it does not compromise the performance of the others. This separation of responsibilities leads to a more reliable overall detection pipeline, reducing the likelihood of false positives and false negatives in critical scenarios [46, 47].
@@
A further significant contribution is the Retrieval-Augmented Generation component, which improves final decision-making by reasoning over retrieved evidence and domain knowledge. This mechanism not only improves classification accuracy, but also generates technically grounded and context-aware explanations. The synergy between a semantic search engine and a summarization-optimized LLM enables the production of comprehensive reports, as confirmed by both automatic metrics (BLEU, ROUGE, BERTScore) and GPT-4-based evaluation.
@@
7 Related Work
The growing interconnection of web services has made the detection and mitigation of sophisticated cyber-attacks increasingly critical. Organizations today face threats ranging from credential theft to large-scale offensives, such as Distributed Denial of Service (DDoS) attacks and Advanced Persistent Threats (APTs). Among the most critical vulnerabilities, web applications represent a prime target for malicious actors who exploit browser flaws, hijack user sessions, or inject harmful code.
Traditional defense mechanisms, based on static rules or known signatures, prove inadequate in the face of rapidly evolving intrusion techniques and the increasing frequency of zero-day attacks. This necessitates the development of more intelligent and adaptive solutions, capable not only of detecting known threats, but also of continuously learning from and adapting to novel attack patterns.
In this context, language models such as BERT [48] offer a powerful opportunity: their ability to understand complex linguistic contexts allows for flexible and generalizable modeling of attack descriptions. Our approach leverages this capability to enable the dynamic integration of more types of attack without requiring a complete system redesign. In table 4 some key difference between classical and AI-powered techniques are proposed.
Aspect Classical Techniques AI-powered Techniques (LLMs)
Adaptability to Novel Attacks Low – limited to known patterns or signatures High – capable of generalization and zero-day inference
Example Tools / Frameworks Firewall, IPS/IDS AutoAttacker, CyberSecEval2, SecBench
Work Focus Area Main Contribution Remarks
Awerbuch et al. [50] MANET and WSN threats Introduces Byzantine-resilient routing to detect and isolate compromised nodes Key reference in secure wireless protocols
Shimeall et al. [52] Cyberterrorism Discusses the strategic and socio-political impact of cyber attacks Highlights need for broader defense frameworks
CyberSecEval 2 [54] Benchmarking LLMs in security Evaluates LLMs on multiple offensive and defensive tasks, including prompt injection Reveals performance/security trade-offs
SecBench [55] Comparative evaluation Provides standardized evaluation metrics for LLM-based security tools Useful for reproducibility and benchmarking
Prompt Injection Defense [57] AI robustness Studies prompt injection vectors and mitigation techniques Highlights vulnerabilities in open LLM interfaces
7.1 Classical techniques
In the context of cyber attack classification and detection systems, “classical” techniques generally focus on the definition of taxonomies, attack patterns, and well-established defensive methods. A traditional aspect involves distinguishing between passive attacks (e.g., eavesdropping and traffic analysis) and active attacks (e.g., injecting malicious packets or tampering with data). Additionally, attacks are often classified into the macro-categories of cyber crime, cyber espionage, cyber terrorism, and cyber war, highlighting how objectives and consequences can vary significantly [59, 60, 61].
An important example of classification is the AVOIDIT taxonomy, which groups attacks according to vectors, objectives, and impacts, underscoring the need for a holistic perspective on security [49]. Within this framework, threats such as Denial of Service (DoS), port scanning, unauthorized data exfiltration, and, more broadly, data integrity violations are analyzed. In parallel, the literature has also focused on distributed attacks (DDoS), where multiple compromised nodes act simultaneously to deny service or cause damage with high impact [62, 63].
In the context of mobile ad hoc networks (MANETs) and wireless sensor networks (WSNs), “classical” classification strategies emphasize the detection of threats such as black hole and wormhole attacks, as well as byzantine variations that compromise nodes internally [50, 64]. Awerbuch et al. [50] offer one of the foundational references to defend against Byzantine attacks in wireless environments, introducing measures aimed at detecting and isolating malicious nodes. Other studies have emphasized the distinction between single-layer and multi-layer attacks, in which adversaries exploit vulnerabilities at various levels of the communication stack [65, 64].
On the prevention and response front, many works focus on the integration of cryptographic tools, authentication systems, and key-management protocols to mitigate the risk of device compromise [66, 63]. This “classical” approach often assumes that the primary defensive perimeter is defined by cryptographic robustness and the proper configuration of network services, prioritizing the use of IDS and IPS [51, 67]. However, as documented in reports on targeted attacks and more complex scenarios such as cyber espionage or cyber terrorism, there is a need for a broader perspective that accounts for socio-technical and strategic aspects [52, 68, 69].
7.2 AI-powered techniques
Recent research underscores the growing impact of AI-based tools, particularly Large Language Models (LLMs), in the domain of cyber-attack orchestration and defense. Early studies demonstrated that LLMs such as ChatGPT could generate attack scripts with success rates ranging from 16% up to 50% when combined with basic cybersecurity skills, thereby lowering the barrier to entry for malicious actors [53, 70]. Offensive platforms like WormGPT and FraudGPT were explicitly developed to harness LLM capabilities for malicious objectives such as phishing, ransomware development, and malware generation [71]. These specialized LLMs highlight a pivotal shift in the cyber-threat landscape, where automation of sophisticated exploits is no longer confined to highly skilled adversaries.
At the same time, the cybersecurity community has begun to leverage LLMs in a defensive capacity, exploiting their ability to interpret contextual nuances in textual data [72, 70]. Advanced frameworks have been proposed to harness LLMs in automated attack classification, attribution, and system hardening. One illustrative example is the AutoAttacker system, which uses LLM-guided modules to orchestrate and evaluate attacks on a simulated organizational network [53]. Meanwhile, works such as CyberSecEval 2 and SecBench emphasize benchmarking and comparative evaluations of LLM-based methods to measure their robustness against real-world offensive scenarios [54, 55].
Despite these advances, multiple studies argue that conventional benchmarks fail to capture the full spectrum of real-world risks posed by LLMs in cybersecurity settings [73]. For instance, emerging methods of prompt injection have exposed how easily an LLM’s responses can be manipulated to generate misleading or harmful content [57]. In response, new lines of research have proposed refining LLMs with domain-specific corpora and advanced retrieval-augmentation techniques (often denoted as RAG: Retrieval-Augmented Generation) to provide more accurate and context-aware threat assessments [56, 74].
Studies focusing on threats at the nation-state level also point to the increasing importance of LLM-driven analysis in attributing sophisticated attacks [58]. Here, combining LLM capabilities with knowledge graphs and advanced threat intelligence fosters deeper insights into malware campaigns and social engineering tactics [75]. Furthermore, reviews of cyber defense LLM applications highlight how these models can enhance intrusion detection, automate forensic analysis, and generate real-time alerts, although persistent challenges remain in interpretability, adversarial robustness, and regulatory compliance [71, 70].
A summary of the works discussed above is provided in Table 5. Within this evolving landscape, CyberRAG draws on the strengths of LLMs to assist in classifying and reporting cyber-attacks while mitigating adversarial manipulations. By adopting a RAG workflow, our approach aims to incorporate an internal knowledge base tailored to cyber threat intelligence, enabling more accurate classification of novel attack vectors and fostering rapid, operator-friendly reporting and mitigation.
@@ 8 Conclusion
This work introduces a modular framework for cyber-attack classification and description, built on the principles of agentic system design and Retrieval-Augmented Generation (RAG). By integrating a specialized Classification tool with a context-aware RAG component, CyberRAG merges the precision of fine-tuned large language models, trained to detect specific web-based malicious payload, with the abstraction capabilities and enriched contextual reasoning offered by the RAG pipeline.
@@ Acknowledgments
This work was partially supported by the SERICS project (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.

### v2

@@ 4.1 Knowledge Bases
For the purposes of this work, we constructed three distinct knowledge bases, one for each type of web-based attack under investigation. The system is designed to distinguish among three specific attack types: SSTI, SQL Injection, and Cross-Site Scripting (XSS). To achieve a deep understanding of these attack vectors, dedicated and categorized documents are required.
@@ 4.2 Attacks’ Dataset
For the Server-Side Template Injection (SSTI) category, no ready-to-use dataset is available. Consequently, we generated this dataset manually using curated payloads from the PayloadAllTheThings SSTI section in conjunction with outputs from the SSTImap tool. The positive class consists of diverse payloads known to trigger SSTI vulnerabilities across multiple template engines. In contrast, the negative class comprises a variety of benign strings such as mail, names or numbers that do not result in SSTI behavior, serving as clean input examples. This dataset can be downloaded by password at [44] an example is provided in Table 4.
payload label
{{{ ”.__class__.__mro__[1].__subclasses__()[59](’/etc/passwd’).read() }}} 1
(3a + 1z) + (9b * 9) = 0 0
Table 4: Extract of SSTI dataset
@@ 5 Experimental Results
More in-depth, each individual component of CyberRAG has been validated, and an assessment of the performance and reliability of the whole system has been performed. This includes the evaluation of the accuracy of the LLM-based classifiers, the effectiveness of the RAG-based explanation module, and the coherence and usability of the final outputs generated by the complete Agentic RAG pipeline.
@@ 5.1 LLM-based Classifier for Attack Identification
Figure 3 shows the classification performance of the BERT-based classifiers (bert-base-uncased, albert-base-v2, distilbert-base-uncased, roberta-base), used to build the classification tool. All models were trained and tested for each attack class considered. Prior to model training, the datasets underwent preprocessing, which involved the removal of incomplete rows, balancing between positive and negative class. Subsequently, each dataset was partitioned into training and test subsets following an 80-20 random split. All classifiers taken into account have been trained with identical hyperparameters: a maximum of 30 epochs and a batch size of 32. The optimizer is RectifiedAdam with a learning rate of $3\times 10^{-5}$. The loss function is binary cross-entropy.
The classifiers were evaluated by using 5 different metrics, three of which are based on the computation of the Area Under the Curve (AUC) for the precision-recall (PR) and receiving operating characteristic (ROC). The other three are the precision (Prec), the F1-score and the binary accuracy (Bin Acc).
@@
From a computational efficiency perspective, each classifier was trained on a NVIDIA A100 GPU with identical hyperparameters. On average across the three attack classes, the training time per model was approximately 2.1 hours, while dataset preparation (tokenization, balancing, and formatting) required less than 10 minutes per class. Once trained, inference proved highly efficient: The average prediction and report generation latency was 0.8 seconds per alert, ensuring the system operates efficiently and meets the time-sensitive demands of enterprise SOC environments.
@@
5.1.1 Model selection rationale and operational constraints
Our evaluation focuses on open-weight encoder models (BERT, RoBERTa, ALBERT, DistilBERT) for two main reasons. First, the artifacts involved in our setting are predominantly textual (payloads, alerts, logs). Encoder architectures remain well suited for fine-grained, bidirectional pattern recognition needed to classify SQLi, XSS, and SSTI, while keeping the training process transparent and reproducible.
Second, open-weight encoders can be fine-tuned on curated security datasets. This capability is essential in a highly specialized and adversarial domain, where domain adaptation materially improves stability and accuracy. By contrast, proprietary frontier models (e.g., GPT-4) do not expose their weights and cannot be fine-tuned; they would need to be used in zero-shot or few-shot configurations only, which typically yields lower and less stable performance for cybersecurity tasks that require targeted adaptation. In addition, closed models entail data egress, cost, and governance constraints that are often incompatible with on-premise, regulated deployments.
@@ 5.3 Evaluation of Generated Explanations
To evaluate the quality of contextualized explanations generated by our RAG framework, we conducted an extensive benchmark across five different open-weight language models, all with approximately 7 billion parameters. The selected models are: DeepSeek-R1 7B [48], Gemma3:4B [49], LLaMA3.1 8B [21], Mistral 7B [50], and Qwen2.5 7B [51]. All of these models were made available and executed within the Ollama environment [52], enabling unified access and deployment for comparative evaluation [^1] At the time the tests were conducted, these represented the state-of-the-art LLMs..
The framework is model-agnostic and forward-compatible: more recent open-weight models (e.g., LLaMA 4, Qwen3) can be integrated as drop-in replacements without structural changes to the system.
@@ 5.3.2 Evaluation via LLM-as-a-Judge
    General Explanation Evaluation: For each generated explanation, GPT-4 rated it on a 1-5 scale for clarity, informativeness, and semantic alignment with the input payload and the attack label.
@@
Table 6: GPT-4 based LLM-judge scores (1-5 scale) per model.
@@ 5.5 Robustness Evaluation
$\text{Correct Classification (\%)}=\frac{\text{\# Correct Predictions}}{\text{\# Predictions}}\times 100\%$ (1)
@@ 6 Discussion
The usage of multiple classifiers allows for a modular architecture that contributes to the robustness of the system. If one classifier performs poorly due to data imbalance or ambiguity in one attack category, it does not compromise the performance of the others. This separation of responsibilities leads to a more reliable overall detection pipeline, reducing the likelihood of false positives and false negatives in critical scenarios.
@@
A further significant contribution is the RAG component, which improves final decision-making by reasoning over retrieved evidence and domain knowledge. This mechanism not only improves classification accuracy, but also generates technically grounded and context-aware explanations. The synergy between a semantic search engine and a summarization-optimized LLM enables the production of comprehensive reports, as confirmed by both automatic metrics (BLEU, ROUGE, BERTScore) and GPT-4-based evaluation.
@@
One of the main limitations of this approach is its dependence on the quality of the data within the knowledge base and the precision of the classifier. A poorly maintained knowledge base or an inaccurate classifier can significantly impact the system’s effectiveness. In addition, the tool is not equipped to describe new types of attacks or attacks it hasn’t been trained to recognize. This limitation makes it less effective against evolving cyber threats, as it can only provide adequate descriptions for attacks that its classifier has been specifically trained on. The effectiveness of CyberRAG is directly dependent on the performance of the preceding IDS/IPS. If the IDS/IPS fails to recognize a security event as potentially dangerous, it won’t generate a log entry for that event. Consequently, CyberRAG will have no data to analyze, rendering it useless in that specific scenario. This highlights a critical vulnerability in the system’s pipeline, where a failure at the initial detection stage prevents any subsequent analysis. We are aware of this limitation. However, given that CyberRAG is designed to operate in contexts with known threats and its effectiveness is dependent on the Knowledge Base, this issue can be mitigated.
It is sufficient to ensure that the rules for the IDS/IPS are constantly updated. To do this, dedicated online services can be used that offer regularly updated rules and signatures, so as to keep the intrusion prevention and detection system always up to date with the most recent and known threats.
Additional studies on detecting threats not based on a priori known signatures are already present in the literature, but they are considered outside the scope of this work.
Building on the approaches mentioned earlier, recent contributions have introduced specialized RAG-based frameworks tailored specifically for cybersecurity tasks. TechniqueRAG [62] addresses the challenge of adversarial technique annotation in low-resource scenarios by combining off-the-shelf retrievers with instruction-tuned LLMs and a novel zero-shot re-ranking strategy. Its main strength lies in improving precision when labeled data are scarce, while CyberRAG focuses on real-time IDS/IPS alert triage and the generation of structured, SOC-ready reports. AURA [63] introduces a multi-agent architecture for the attribution of APTs, integrating heterogeneous sources such as Trusted Third Parties (TTPs), Indicators of Compromise (IoCs), and malware artifacts. In contrast, CyberRAG targets a different stage of the defense pipeline, aiming to reduce false positives and provide interpretable explanations at the alert triage level. Another complementary direction is CTIKG [64], which enriches RAG pipelines with knowledge graph structures to enhance cyber threat intelligence and contextual reasoning.
@@ 7 Conclusion
This work introduces a modular framework for cyber-attack classification and description by integrating a specialized Classification tool with a context-aware RAG component. CyberRAG merges the precision of fine-tuned large language models, trained to detect specific web-based malicious payload, with the abstraction capabilities and enriched contextual reasoning offered by the RAG pipeline.
@@ Acknowledgments
This work was partially supported by the projects SERICS (PE00000014) and FAIR (PE0000013) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
@@
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to check for grammar errors, typos, and overall writing clarity. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.